Two trains leave a city at the same time. One travels north, and the other travels south 20 mph. faster. In 2 hr, the trains are 280 miles apart. Find their speeds.
Answer by checkley71(8403)

2x+2(x+20)=280 or 2x+2x+40=280 or 4x=280-40 or 4x=240 x=240/4 or x=60 mph.